# AudioGenie: A Training-Free Multi-Agent Framework for Diverse Multimodality-to-Multiaudio Generation

### v1

@@ Abstract.
Multimodality-to-Multiaudio (MM2MA) generation faces significant challenges in synthesizing diverse and contextually aligned audio types (e.g., sound effects, speech, music, and songs) from multimodal inputs (e.g., video, text, images), owing to the scarcity of high-quality paired datasets and the lack of robust multi-task learning frameworks. Recently, multi-agent system shows great potential in tackling the above issues. However, directly applying it to MM2MA task presents three critical challenges: (1) inadequate fine-grained understanding of multimodal inputs (especially for video), (2) the inability of single models to handle diverse audio events, and (3) the absence of self-correction mechanisms for reliable outputs. To this end, we propose AudioGenie, a novel training-free multi-agent system featuring a dual-layer architecture with a generation team and a supervisor team. For the generation team, a fine-grained task decomposition and an adaptive Mixture-of-Experts (MoE) collaborative entity are designed for dynamic model selection, and a trial-and-error iterative refinement module is designed for self-correction. The supervisor team ensures temporal-spatial consistency and verifies outputs through feedback loops. Moreover, we build MA-Bench, the first benchmark for MM2MA tasks, comprising 198 annotated videos with multi-type audios. Experiments demonstrate that our AudioGenie outperforms state-of-the-art (SOTA) methods across 9 metrics in 8 tasks. User study further validate the effectiveness of the proposed method in terms of quality, accuracy, alignment, and aesthetic. The anonymous project website with samples can be found at https://audiogenie.github.io/.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Collaborative and social computingccs: Information systems Multimedia content creation
@@ 1. Introduction
Recently, autoagent-based paradigms have shown potential in handling multi-task (Chen et al., 2024a; Huang et al., 2024; Liang et al., 2024) and data-scarce scenarios (Li et al., 2024b; Tu et al., 2024). However, directly applying the vanilla autoagent-based paradigm to MM2MA generation suffers from three main challenges: (1) Input Understanding Issue: Limited ability of understanding fine-grained multimodal inputs. Multimodal inputs, especially video, require fine-grained understanding and temporal-spatial consistency capturing, which existing vanilla agent methods struggle to handle. (2) Selecting Issue: A single agent or model cannot effectively manage all types of audio events. Each audio type requires domain-specific expertise. Moreover, even within the same task, different models often posses different strengths and focus. It is nearly impossible for one well-trained model to satisfy all user needs, and relying on a single model is typically suboptimal. (3) Self-Correcting Issue: Models lack the ability to automatically self-correct results, leading to unreliable outputs. Existing pipelines typically generate audio in a single pass without iterative feedback loops, making it hard to fix errors or integrate late-stage adjustments. Even well-trained models may misalign audio events with user inputs or produce artifacts that remain unchecked. In fact, some agent-based methods (e.g.,  (Zhang et al., 2025b)) have been explored for single-input single-output task (e.g., video to sound-effect generation), yet they still face challenges mentioned above.
@@ 2.1. X-to-Audio Generation
X-to-audio generation, which aims to synthesize various types of audio such as music, sound effects, speech, and song from various inputs (i.e., $X$ modality can be video, text, images or the combination of them), is an emerging area of research that faces significant challenges in multi-modal understanding and generation.
Existing audio generation models often focus on narrow input modalities or restricted conditioning types. For instance, text-to-audio approaches focus on generating audio signals from textual descriptions or instructions, like speech (Du et al., 2024; Guo et al., 2024a; Zhou et al., 2024; Xie et al., 2024a; Rong and Liu, 2025; Rong et al., 2025; Chen et al., 2024b), sound effect (Yuan et al., 2024; Huang et al., 2023; Liu et al., 2023; Xue et al., 2024), music (Copet et al., 2023; Chen et al., 2024c; Melechovsky et al., 2023; Liu et al., 2024b), or song (Lei et al., 2024b; Hong et al., 2024; Ding et al., 2024; Liu et al., 2025). While video-to-audio methods focusing on generating either speech (Yemini et al., 2023; Lei et al., 2024a; Choi et al., 2025; Gupta et al., 2024; Kim et al., 2025), music (Zhuo et al., 2023; Li et al., 2024a, c; Xie et al., 2025a; Zuo et al., 2025), or sound effects (Luo et al., 2023; Xie et al., 2024b; Wang et al., 2024c, a; Comunità et al., 2024; Ren et al., 2025; Zhang et al., 2025b), which are synchronized to visual cues. These models excel in specialized tasks, but limit their adaptability to complex multi-modal scenarios.
The growing demand for holistic audio synthesis has spurred interest in unified frameworks that can generate multiple types of audio with multi-modal condition. For example, AudioX (Tian et al., 2025) proposes a unified model capable of flexibly handling various inputs, such as text, video, images, music, and audio, to generate general-purpose audio and music. However, training these models requires large-scale video-audio pairs, leading to the creation of two comprehensive datasets for training.
Instead of relying on extensive training, we propose a training-free unified multi-agent system that achieves efficient collaborative generation between multi-modal inputs and multiple types of audio outputs.
@@ 2.2. Auto-agent System
Several studies utilize multimodal large language models (MLLMs) as core AI agents, enabling intelligent planning and decision-making to address more complex tasks, such as image generation and editing (Wang et al., 2024b), video generation (Tu et al., 2024; Li et al., 2024b), tool usage (Qu et al., 2025; Gao et al., 2025), and system development (Qian and Cong, 2023).
Although MLLM agents have been applied to many tasks, their use in audio generation remains limited. Some methods leverage agents to perform multiple audio-related tasks. For example, AudioGPT (Huang et al., 2024) and WavCraft (Liang et al., 2024) directly connect multiple audio foundation models for speech, music, and sound understanding and simple generation. Recently, LVAS-Agent (Zhang et al., 2025b) proposes a multi-agent framework that emulates professional dubbing workflows through collaborative role specialization for long-video synthesis. However, as we mentioned above, it mainly concentrates on sound effects and music, and its training process is still computationally intensive.
@@ 3.1. AudioGenie: Dual-Layer Coordinated Structure
Single-pass generation methods often struggle to produce high-quality outputs for complex and multi-event audio requirements. To address this, we propose the AudioGenie with a Dual-Layer Coordinated Structure, consisting of a generation team and a supervisor team.
@@
Supervisor Team. Prior work has shown that assigning complex reasoning tasks to specialized sub-roles can improve both efficiency and quality (Zhang et al., 2023). Considering this, the supervisor team monitors and provides feedback at each stage through a dynamic domain-specific expert simulation way, where supervisor is role-conditioned to focus on a specific aspect of the workflow. Specifically, in the first stage, the supervisor acts as an audio planning expert to evaluate the plan from content suitability, timing accuracy, and alignment, and provide concrete suggestions and revise them if necessary. In the second stage, the supervisor acts as a domain expert to assess model choices and refined JSON files for each audio event type. In the third stage, supervisor serves as a professional audio evaluator, collaborating with the generation team to perform on-the-fly self-corrections.
This coordinated “plan-and-verify” paradigm fosters effective collaboration among multiple agents, maintaining flexibility while ensuring that final outputs remain closely aligned with input requirements.
@@ 3.2. Fine-Grained Task Decomposition Module
Multimodal inputs often contain multiple overlapping or interacting long-duration audio events. To achieve fine-grained coordination across time, content, and style for these events, we design a fine-grained task decomposition module that splits the overall audio requirements into a series of manageable sub-audio events. Specifically, the agent analyzes the semantic and temporal information in the input to identify implied audio events. Each aduio event is assigned an audio type (i.e., speech, sound effect, song, or music), a description, and a timestamp. The resulting plan is represented as a structured JSON list with content {audio_type, start_time, end_time, duration, description, volume}. For any included visual information, the agent is prompted to generate captions that describe the overall scene, providing a unified ambience for subsequent style planning.
@@
Video-to-Sound Effect Generation MMAudio (Cheng et al., 2024)
Text-to-Sound Effect Generation MMAudio (Cheng et al., 2024)
@@ 3.4. Trial-and-Error Iterative Refinement Module
Single-pass audio generation often suffers from unstable output quality and poor coherence. Inspired by  (Wang et al., 2024b), we propose the Trial-and-Error Iterative Refinement Module. We organize the generation process of each audio event as a tree-of-thought structure, enabling the agent to explore multiple generation paths and adjust or backtrack as needed. This structure consists of three types of nodes: an initial node, generation nodes, and refinement nodes. The initial node is the root of the tree and receives the corresponding JSON file, marking the beginning of the generation. The generation node produces the audio event based on the planned task by using the candidate base generation model. The refinement node is built based on feedback from the supervisor team to adjust the specific quality issues or defects.
@@
By exploring different paths during generation, our module retains flexibility: if one branch fails to produce a satisfactory result, the system can backtrack and try an alternate path. The use of refinement nodes and retry sibling generation nodes embodies the trial-and-error paradigm, as the system continuously evaluates the results, applies corrections or alternatives, and iteratively refines the output.
@@ 4. Experiment
4.1. Dataset
We construct a specialized Multi-type Audio synthesis Benchmark (MA-Bench) comprising 198 professionally curated videos sourced from three main origins: (1) short videos from YouTube or Bilibili; (2) video material websites; and (3) synthetic scenes generated by Open Sora  [^1] https://openai.com/sora/. Importantly, every video in the dataset features multi-type audio tracks, including diverse sound effects, background music, and human speech or singing. To ensure diversity, VABench spans a wide range of video categories. Figure 2 (a) quantifies the distribution across 7 video-level categories, while Figure 2 (b) shows video frames and its corresponding annotations from three data source respectively.
4.2. Implementation Details
@@ 4.4. Comparison with SOTA Methods
In the text-input and image-input settings, we compare our method with four text-to-audio generation approaches (Auffusion (Xue et al., 2024), MMAudio (Cheng et al., 2024), Make-An-Audio (Huang et al., 2023), and Tango 2 (Majumder et al., 2024)) and two image-to-audio approaches (See2Sound (Dagli et al., 2024) and SSV2A (Guo et al., 2024b)), respectively. Our method achieves significantly higher performance on multiple metrics. Existing methods struggle with complex text instructions or images containing multiple audio events, often yielding incomplete or inaccurate generations. Moreover, their outputs commonly exhibit style mismatches relative to the scene in the input image or the theme of the input text.
@@
Make-An-Audio (Huang et al., 2023) 0.478 5.498 2.553 2.998 4.659 $1.67\pm 0.50$ $1.22\pm 0.44$ $1.11\pm 0.33$ $1.22\pm 0.44$
@@ 4.5. Ablation Studies
We further examine the importance of different modules through the following ablation settings: (1) w/o supervisor team: remove the supervisor team in the dual-layer coordinated structure, leaving only the generation team; (2) w/o domain-specific experts: replace the domain-specific experts in stage 2 with a single general expert, who handles script refinement and model selection for all audio types; (3) w/o intra-expert self-refinement: disable this mechanism; (4) w/o inter-expert collaborative refinement: disable this mechanism; and (5) w/o Tree-of-thought self-correction: disable trial-and-error iterative refinement in stage 3. The results are shown in Table 5.
@@ 5. Conclusion
In this work, we propose a novel training-free multi-agent system, AudioGenie, which supports multimodal input and multiple types of audio outputs. We introduce a dual-layer coordinated structure comprising a generation team and a supervisor team, which handle the planning and execution of generation, as well as the supervision and correction of each step, respectively. For the generation team, a fine-grained task decomposition and an adaptive Mixture-of-Experts collaborative entity are designed for detailed multimodal input understanding and dynamic model selection. We also propose a trial-and-error iterative refinement module that constructs a tree-of-thought for each audio event, enabling continuous improvements by evaluating and revising intermediate outcomes. In addition, we build MA-Bench, the first benchmark for MM2MA generation, which includes 198 annotated videos with multiple audio types. Experimental results demonstrate that our approach produces complex multi-audio events aligned with multimodal inputs, surpassing SOTA methods across several metrics. Qualitative assessments and user studies further confirm the effectiveness of our model in terms of quality, alignment, and aesthetics.
@@ Appendix A Additional SOTA Comparisons on Short Videos
Given that FoleyCrafter (Zhang et al., 2024) only supports audio under 10 seconds, we further evaluated different methods on short videos for a more comprehensive comparison, considering both video-only and video+text inputs. The results are shown in Table 6. Specifically, for video-only inputs, our approach is slightly lower on the objective metrics PAM and PQ compared to other methods, but it achieves the best performance on all subjective metrics, and outperforms others on PC, CE, and CU. In the video+text input scenario (with the same experimental setting in Table 3), our method attains the best results in terms of audio quality, completeness, alignment, and aesthetics.
Table 6. Comparison with SOTA methods. For a fair comparison with FoleyCrafter, we select videos shorter than 10 seconds. Best performances are highlighted in bold, while second-best are underlined.
Input Modalities Methods Objective Metrics Subjective Metrics
Video MMAudio (Cheng et al., 2024) 0.645 6.519 2.930 3.881 6.224 $\underline{3.09}\pm 1.33$ $\underline{2.70}\pm 1.20$ $\underline{2.36}\pm 1.34$ $2.30\pm 1.06$
FoleyCrafter (Zhang et al., 2024) 0.568 6.255 2.891 3.420 5.928 $2.67\pm 1.42$ $2.44\pm 1.35$ $2.22\pm 1.25$ $\underline{2.33}\pm 0.97$
Seeing and Hearing (Xing et al., 2024) 0.337 5.142 3.034 2.936 4.760 $2.10\pm 1.20$ $2.30\pm 1.25$ $2.09\pm 1.25$ $1.80\pm 0.92$
AudioX (Tian et al., 2025) 0.437 5.656 2.915 3.340 5.194 $2.37\pm 1.34$ $2.19\pm 1.55$ $2.27\pm 1.75$ $1.91\pm 1.10$
Ours 0.558 6.448 4.092 4.730 6.239 $\textbf{3.90}\pm 0.99$ $\textbf{3.50}\pm 0.85$ $\textbf{3.88}\pm 0.84$ $\textbf{3.70}\pm 0.94$
Combination (Video + Text) MMAudio (Cheng et al., 2024) 0.841 5.935 4.573 5.098 5.797 $\underline{2.50}\pm 1.37$ $\underline{1.67}\pm 0.81$ $1.33\pm 0.51$ $\underline{2.17}\pm 1.47$
FoleyCrafter (Zhang et al., 2024) 0.391 6.073 4.448 3.476 5.537 $1.50\pm 0.54$ $1.28\pm 0.48$ $1.13\pm 0.35$ $1.33\pm 0.52$
AudioX (Tian et al., 2025) 0.425 5.806 3.859 4.563 4.988 $1.71\pm 0.75$ $1.43\pm 0.53$ $\underline{1.57}\pm 0.78$ $1.28\pm 0.75$
Ours 0.651 6.465 5.261 5.150 5.223 $\textbf{4.86}\pm 0.38$ $\textbf{4.67}\pm 0.51$ $\textbf{4.71}\pm 0.48$ $\textbf{4.50}\pm 0.55$

### v2

@@ Abstract.
Multimodality-to-Multiaudio (MM2MA) generation faces significant challenges in synthesizing diverse and contextually aligned audio types (e.g., sound effects, speech, music, and songs) from multimodal inputs (e.g., video, text, images), owing to the scarcity of high-quality paired datasets and the lack of robust multi-task learning frameworks. Recently, multi-agent system shows great potential in tackling the above issues. However, directly applying it to MM2MA task presents three critical challenges: (1) inadequate fine-grained understanding of multimodal inputs (especially for video), (2) the inability of single models to handle diverse audio events, and (3) the absence of self-correction mechanisms for reliable outputs. To this end, we propose AudioGenie, a novel training-free multi-agent system featuring a dual-layer architecture with a generation team and a supervisor team. For the generation team, a fine-grained task decomposition and an adaptive Mixture-of-Experts (MoE) collaborative entity are designed for detailed comprehensive multimodal understanding and dynamic model selection, and a trial-and-error iterative refinement module is designed for self-correction. The supervisor team ensures temporal-spatial consistency and verifies outputs through feedback loops. Moreover, we build MA-Bench, the first benchmark for MM2MA tasks, comprising 198 annotated videos with multi-type audios. Experiments demonstrate that our AudioGenie achieves state-of-the-art (SOTA) or comparable performance across 9 metrics in 8 tasks. User study further validates the effectiveness of our method in terms of quality, accuracy, alignment, and aesthetic. The project website with audio samples can be found at https://audiogenie.github.io/.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755758isbn: 979-8-4007-2035-2/2025/10ccs: Information systems Multimedia content creation
@@ 1. Introduction
Recently, autoagent-based paradigms have shown potential in handling multi-task (Chen et al., 2024a; Huang et al., 2024; Liang et al., 2024) and data-scarce scenarios (Li et al., 2024a; Tu et al., 2024). However, directly applying the vanilla autoagent-based paradigm to MM2MA generation suffers from three main challenges: (1) Input Understanding Issue: Limited ability of understanding fine-grained multimodal inputs. Multimodal inputs, especially video, require fine-grained understanding and temporal-spatial consistency capturing, which existing vanilla agent methods struggle to handle. (2) Selecting Issue: A single agent or model cannot effectively manage all types of audio events. Each audio type requires domain-specific expertise. Moreover, even within the same task, different models often posses different strengths and focus. It is nearly impossible for one well-trained model to satisfy all user needs, and relying on a single model is typically suboptimal. (3) Self-Correcting Issue: Models lack the ability to automatically self-correct results, leading to unreliable outputs. Existing pipelines typically generate audio in a single pass without iterative feedback loops, making it hard to fix errors or integrate late-stage adjustments. Even well-trained models may misalign audio events with user inputs or produce artifacts that remain unchecked. In fact, some agent-based methods (e.g.,  (Zhang et al., 2025b)) have been explored for single-input single-output task (e.g., video to sound-effect generation), yet they still face challenges mentioned above.
@@ 2.1. X-to-Audio Generation
X-to-audio generation, which aims to synthesize various types of audio such as music, sound effects, speech, and song from various inputs (i.e., $X$ modality can be video, text, images or any combination), is an emerging area of research that faces significant challenges in multi-modal understanding and generation.
Current audio generation models typically target a single input modality or a limited conditioning type. For instance, text-to-audio methods generate speech (Du et al., 2024; Guo et al., 2024a; Xie et al., 2024; Rong and Liu, 2025; Rong et al., 2025), sound effects (Yuan et al., 2024; Huang et al., 2023; Liu et al., 2023; Xue et al., 2024), music (Copet et al., 2023; Chen et al., 2024c; Melechovsky et al., 2023; Liu et al., 2024b), or songs (Lei et al., 2024b; Hong et al., 2024; Ding et al., 2024; Liu et al., 2025) from textual descriptions. Video-to-audio approaches produce speech (Lei et al., 2024a; Choi et al., 2025; Gupta et al., 2024; Kim et al., 2025), music (Tian et al., 2024; Li et al., 2024b; Xie et al., 2025a; Zuo et al., 2025; Wang et al., 2025), or sound effects (Cheng et al., 2024; Wang et al., 2024b; Comunità et al., 2024; Ren et al., 2025; Zhang et al., 2025b) synchronized with visual cues. While these specialized models excel in their respective tasks, they lack flexibility for more complex multimodal scenarios.
The demand for holistic audio synthesis has driven the development of unified frameworks capable of generating multiple audio types under multimodal conditions. For instance, AudioX (Tian et al., 2025) presents a single model that flexibly handles text, video, and image inputs to produce both sound effects and music. However, such models are data-hungry, demanding large-scale paired datasets and costly training procedures.
Instead of relying on extensive training, we propose a training-free unified multi-agent system for efficient, collaborative generation of multiple audio types from multimodal inputs.
@@ 2.2. Auto-agent System
Several studies utilize multimodal large language models (MLLMs) as core AI agents, enabling intelligent planning and decision-making to address more complex tasks, such as image generation and editing (Wang et al., 2024a), video generation (Tu et al., 2024; Li et al., 2024a), and system development (Qian and Cong, 2023).
Although MLLM agents have been applied to many tasks, their use in audio generation remains limited. Early systems like AudioGPT (Huang et al., 2024) and WavCraft (Liang et al., 2024) use an agent to coordinate multiple expert audio models for basic understanding and generation. More recently, LVAS-Agent (Zhang et al., 2025b) proposes a multi-agent framework for long-video audio synthesis. However, this method focuses only on sound effects.
@@ 3.1. AudioGenie: Dual-Layer Coordinated Structure
Single-pass generation methods often struggle to produce high-quality audio for complex, multi-event scenarios. To address this, we introduce AudioGenie with a dual-layer coordinated structure, composed of a generation team and a supervisor team.
@@
Supervisor Team. Prior work has shown that assigning complex reasoning tasks to specialized roles can improve both efficiency and quality (Zhang et al., 2023). Considering this, the supervisor team monitors and provides feedback at each stage through a dynamic domain-specific expert simulation way, where supervisor is role-conditioned to focus on a specific aspect. Specifically, in the first stage, the supervisor acts as an audio planning expert to evaluate the plan from content suitability, timing accuracy, and alignment, and provide concrete suggestions and revise them if necessary. In the second stage, the supervisor acts as a domain expert to assess model choices and refined JSON files for each audio event type. In the third stage, supervisor serves as a professional audio evaluator, collaborating with the generation team to perform on-the-fly self-corrections.
This coordinated “plan-and-verify” paradigm fosters effective collaboration among multiple agents, maintaining flexibility while ensuring that final outputs remain aligned with input requirements.
@@ 3.2. Fine-Grained Task Decomposition Module
Multimodal inputs often contain multiple overlapping or interacting long-duration audio events. To achieve fine-grained coordination across time, content, and style for these events, we design a fine-grained task decomposition module that splits the overall audio requirements into a series of manageable sub-audio events. Specifically, the agent analyzes the semantic and temporal information in the input to identify implied audio events. Each aduio event is assigned an audio type (i.e., speech, sound effect, song, or music), a description, and a timestamp. The resulting plan is represented as a structured JSON list with content {audio_type, object, start_time, end_time, description, volume}. For any included visual information, the agent is prompted to generate captions that describe the overall scene, providing a unified ambience for subsequent style planning.
@@
Video/Text-to-Sound Effect Generation MMAudio (Cheng et al., 2024)
@@ 3.4. Trial-and-Error Iterative Refinement Module
Single-pass audio generation often suffers from unstable output quality and poor coherence. Inspired by  (Wang et al., 2024a), we propose the Trial-and-Error Iterative Refinement Module. We organize the generation process of each audio event as a tree-of-thought structure, enabling the agent to explore multiple generation paths and adjust or backtrack as needed. This structure consists of three types of nodes: an initial node, generation nodes, and refinement nodes. The initial node is the root of the tree and receives the corresponding JSON file, marking the beginning of the generation. The generation node produces the audio event based on the planned task by using the candidate base generation model. The refinement node is built based on feedback from the supervisor team to adjust the specific quality issues or defects.
@@
By exploring different paths during generation, our module retains flexibility: if one branch fails to produce a satisfactory result, the system can backtrack and try an alternate path. The use of refinement nodes and retry sibling generation nodes embodies the trial-and-error paradigm, which facilitate a continuous loop of evaluation, correction, and iterative improvement of the output.
@@ 4. Experiment
4.1. Dataset
We introduce the Multi-type Audio synthesis Benchmark (MA-Bench), comprising 198 professionally curated videos sourced from three main origins: (1) online platforms (YouTube, Bilibili); (2) video material websites; and (3) synthetic generators (Sora (Brooks et al., 2024)). Importantly, every video in the dataset features multi-type audio tracks, including diverse sound effects, background music, and human speech or singing. To ensure diversity, MABench spans a wide range of video categories. Figure 2 (a) quantifies the distribution across 7 video-level categories and Figure 2 (b) shows video frames and its corresponding annotations from three data sources.
4.2. Implementation Details
@@ 4.4. Comparison with SOTA Methods
In the text-input and image-input settings, we compare our method with five text-to-audio generation approaches (MMAudio (Cheng et al., 2024), Auffusion (Xue et al., 2024), AudioX (Tian et al., 2025), Make-An-Audio (Huang et al., 2023), and Tango 2 (Majumder et al., 2024)) and two image-to-audio approaches (See2Sound (Dagli et al., 2024) and SSV2A (Guo et al., 2024b)), respectively. Our method achieves significantly higher performance on multiple metrics. Existing methods struggle with complex text instructions or images containing multiple audio events, often yielding incomplete or inaccurate generations. Moreover, their outputs commonly exhibit style mismatches relative to the scene in the input image or the theme of the input text.
@@
Make-An-Audio (Huang et al., 2023) 0.478 5.498 2.553 2.998 4.659 $1.67\pm 0.50$ $22\pm 0.44 1.1$$11\pm 0.33$ $1.22\pm 0.44$
@@ 4.5. Ablation Studies
We further examine the importance of different modules through the following ablation settings: (1) w/o supervisor team: remove the supervisor team in the framework, leaving only the generation team; (2) w/o domain-specific experts: replace the domain-specific experts in stage 2 with a single general expert, who handles script refinement and model selection for all audio types; (3) w/o intra-expert self-refinement: disable this mechanism; (4) w/o inter-expert collaborative refinement: disable this mechanism; and (5) w/o Tree-of-thought self-correction: disable trial-and-error iterative refinement in stage 3. The results are shown in Table 5.
@@ 5. Conclusion
In this work, we propose a novel training-free multi-agent system, AudioGenie, which supports multimodal input and multiple types of audio outputs. We introduce a dual-layer coordinated structure comprising a generation team and a supervisor team, which handle the planning and execution of generation, as well as the supervision and correction of each step, respectively. For the generation team, a fine-grained task decomposition and an adaptive MoE collaborative entity are designed for detailed multimodal input understanding and dynamic model selection. We also propose a trial-and-error iterative refinement module that constructs a tree-of-thought for each audio event, enabling continuous improvements by evaluating and revising intermediate outcomes. In addition, we build MA-Bench, the first benchmark for MM2MA generation, which includes 198 annotated videos with multiple audio types. Experiments demonstrate that our AudioGenie achieves SOTA or comparable performance across several metrics. Qualitative assessments and user studies further confirm the effectiveness of our model in terms of quality, accuracy, alignment, and aesthetic.
@@ Appendix A Additional SOTA Comparisons on Short Videos
Given that FoleyCrafter (Zhang et al., 2024) only supports audio under 10 seconds, we further evaluate different methods on short videos for a more comprehensive comparison, considering both video-only and video+text inputs. The results are shown in Table 6. Specifically, for video-only inputs, our approach is slightly lower on the objective metrics PAM and PQ compared to other methods, but it achieves the best performance on all subjective metrics, and outperforms others on PC, CE, and CU. In the video+text input scenario (with the same experimental setting in Table 3), our method attains the best results in terms of audio quality, completeness, alignment, and aesthetics.
Appendix B Generalization of results to public datasets
We further evaluate AudioGenie on two widely used public datasets, VGGSound (Chen et al., 2020) and V2M (Tian et al., 2024), with quantitative results in Table 7. For single-event generation, our method achieves performance comparable to SOTA V2A models. However, in complex, multi-event scenarios, existing methods often collapse to a single dominant event, leading to incomplete audio generation. In contrast, our method successfully captures multiple events, ensuring better spatio-temporal and semantic alignment with the input video. Furthermore, the audio generated by our framework exhibits superior quality and realism.
Appendix C Results on Different MLLMs
To show the scalability of our framework, we replace Qwen2.5-VL-72B-Instruct with GPT-4o (OpenAI, 2024) and Gemini2.5-Pro (DeepMind, 2025) and tested on the V2A task. As detailed in Table 7, our method outperforms existing SOTA V2A models when using any of these MLLMs as the agent. We also observe that different MLLMs exhibit distinct advantages on various metrics, indicating that our framework is not reliant on a specific MLLM.